OpenStudy (anonymous):

Bob has two savings accounts. He deposited $150 more into account B than account A. After a period of time, account A has earned $30 in interest at 3% and account B has earned $70 in interest at a rate of 4%. 1.Write an equation to represent the situation. Explain each variable used. . 2.How much money did Bob initially deposit into each account? Solve the equation.

OpenStudy (paxpolaris):

let a be the money initially deposited in A b be the money initially deposited in B t be the time

OpenStudy (paxpolaris):

the 3 equations you get are: \[\large b= a+150\] and from i=p*r*t : \[\large 70 = b \times 0.04 \times t\]\[\large 30 = a \times 0.03 \times t\]

OpenStudy (anonymous):

what is the i=p*r*t there for

OpenStudy (paxpolaris):

the general equation for simple interest that u use to get the other 2 equations

OpenStudy (anonymous):

ok thx and now part 2 plzzz

OpenStudy (paxpolaris):

not sure if all of that should be represented by a single equation... for part 2. divide the 2nd equation by the 3rd equation \[\large {70 \over 30} = {b \times0.04\times t \over a \times0.03\times t }\]

OpenStudy (paxpolaris):

\[\large \implies \frac 73= {4b \over 3 a}\]\[\large \implies 7a=4b\]\[\large \implies 7a=4\left( a+150 \right)\]

OpenStudy (paxpolaris):

you can solve that equation ... to find a.

OpenStudy (anonymous):

yes thx soooooooooooo much....if i send u another queation in a message will u answer it for me too

OpenStudy (paxpolaris):

a=$200 b=$350 close this question and ask a new one... i'll see if i can answer

OpenStudy (anonymous):

ok thx sooo much